# A Random Rule Model

### 04 Mar 2026 v1

@@ Appendix B Proof of Theorem 1
By (G1), $\mathrm{rank}(H^{(k)})=|\mathcal{F}|-1$, so the nullspace of $H^{(k)}$ is one-dimensional: there exists a nonzero vector $\bar{\omega}^{(k)}$ such that every solution to $H^{(k)}\omega=0$ is of the form $\omega=c\,\bar{\omega}^{(k)}$ for some scalar $c$. Because the model weights satisfy $\omega_{f}(x)>0$, the relevant solutions have $c>0$. Thus, the menus with $\psi(A)=x^{(k)}$ identify $\omega(x^{(k)})$ up to a positive multiplicative scale factor.
@@
Appendix C Attention-Rule Removal
@@
Figures A.1 and A.2 replicate the complexity and risk mechanism plots (Figures 6 and 7) for MMn, MMx, MAP, SAL under the 10-rule specification (no A1/A2), compared with the full 12-rule library. The qualitative patterns persist: MMx rises with tradeoff complexity while MAP declines, and along risk asymmetry, extreme-outcome comparisons decrease while MAP increases. The removal of the attention rules reallocates their former responsibility mass to the structured rules, but the relative ordering and directional movements of the non-attention rules remain intact. This confirms that the interpretability claims from the mechanism analysis concern genuine behavioral regularities in the structured rules, not patterns induced by the presence of the attention channel.
@@
Appendix D Additional Figures
Figure A.3: Mean effective usage weights across deciles of tradeoff complexity $\mathrm{Comp}(A)$ for all rules.
Figure A.4: Mean effective usage weights across deciles of risk asymmetry $\mathrm{RiskAsym}(A)$ for all rules.
Appendix E Paired Comparisons
@@
Context-dependent Rule-gating test_mse -0.00120 0.00011 50
Neural CPT Rule-gating test_mse 0.00681 0.00013 50
Neural EU Rule-gating test_mse 0.00751 0.00012 50
Neural PT Rule-gating test_mse 0.00599 0.00013 50
Rule-gating Rule-gating test_mse -0.00183 0.00006 50
Value-based Rule-gating test_mse 0.00457 0.00012 50
MOT Rule-gating test_mse -0.00325 0.00011 50
Appendix F Model-Level Uncertainty Intervals
@@
Appendix G Ablation Effect Split-Variability Intervals
@@
The near-identical ablation effects for MMa, REG, and REGmed are not a computational artifact: these three rules have very small and noisy effects because they are largely substitutable with the other rules in the library. Their ablation impacts are indistinguishable at the reported precision.
@@
Appendix H Decomposition Stability
@@
A2 0.2333 0.0118 0.05
A1 0.2057 0.0121 0.06
MMx 0.1214 0.0067 0.06 0.0557 0.0278 0.50
SAL 0.1059 0.0045 0.04 0.0626 0.0305 0.49
REG 0.0628 0.0049 0.08 0.0207 0.0262 1.26
MAP 0.0590 0.0068 0.11 0.0112 0.0195 1.74
MMn 0.0574 0.0225 0.39 0.0266 0.0220 0.83
REGmed 0.0508 0.0110 0.22 0.0255 0.0241 0.95
SAL2 0.0416 0.0023 0.05 0.0794 0.0372 0.47
DISmed 0.0296 0.0070 0.24 0.0125 0.0192 1.54
DIS 0.0237 0.0100 0.42 0.0144 0.0194 1.35
MMa 0.0088 0.0039 0.45 0.0029 0.0158 5.47
Note: Spearman rank correlation of $w_{f}$ orderings across all fold pairs: mean $\rho=0.935$, range $[0.776,1.000]$. Spearman rank correlation of $\phi(f)$ orderings: mean $\rho=0.564$, range [0.794,0.976][-0.794,0.976].
Appendix I Selection Stability
@@
Appendix J Menu-Level Portability (Legacy)
Table A.7 reports the legacy menu-level portability metrics for comparability with the prior version.
Table A.7: Portability to CPC18: out-of-domain MSE when models are trained on choices13k and evaluated on CPC18 without re-tuning. Lower is better.
Model $\mathrm{MSE}_{\mathrm{CPC}}$ $\mathrm{MSE}_{\mathrm{CPC}}$ (trial-wtd) # menus # trials
Rule-gating 0.0170 0.0172 180 354,400
Context-dependent 0.0172 0.0178 180 354,400
Neural CPT 0.0320 0.0313 180 354,400
Neural EU 0.0320 0.0316 180 354,400
Value-based 0.0321 0.0321 180 354,400
Neural PT 0.0326 0.0322 180 354,400
MOT 0.0926 0.0892 180 354,400
Appendix K Feature-Based Baselines: Details
We include six feature-based baselines as additional benchmarks. Each uses the same 50-fold CV protocol and evaluation metrics as the main models.
Feature sets.
Three feature sets are used:
  • F1: Summary statistics ($\approx$18 dimensions): expected value, variance, standard deviation, min, max, range, and number of support points for each lottery, plus pairwise gaps (EV gap, variance gap, min gap, max gap).
  • F2: Rule features (24 dimensions): for each menu $A$ and each rule $f\in\mathcal{F}$, we concatenate the left-recommendation indicator $L_{f}(A)$ and the activity indicator $A_{f}(A)$.
  • F3: Combined ($\approx$42 dimensions): concatenation of F1 and F2.
Models and tuning.
  • Fractional logit: we fit a binomial GLM with logit link (statsmodels.GLM, family=Binomial()) directly on the continuous choice frequency $\hat{p}(A)\in[0,1]$, following Papke and Wooldridge (1996). The $L_{2}$ regularization parameter $\alpha$ is selected from $\{0,0.01,0.1,1\}$ by inner-fold validation (80/20 split within training).
  • Gradient-boosted trees (GBT): we fit GradientBoostingRegressor (scikit-learn) with max_depth=4. The number of estimators is selected from $\{100,200,500\}$ by inner-fold validation. The target is the continuous choice frequency $\hat{p}(A)$.
All hyperparameter selection is restricted to training folds; no CPC18 or test-set information is used.
Appendix L Placebo Rule Library
A distinct concern from library selection is whether the rule indicators themselves carry genuine predictive content, or whether the gate could achieve similar performance with any set of binary indicators that have the same marginal statistics. To address this, we construct a placebo test in which the rule indicators are scrambled while preserving their marginal properties.
Stratified permutation design.
A naïve permutation of rule indicators across all menus would change the marginal activity and recommendation rates in ways that could trivially degrade performance. To avoid this confound, we stratify menus by a simple measure of menu complexity-the total number of support points across the two lotteries, $|\mathrm{Supp}(L^{1})|+|\mathrm{Supp}(L^{2})|$-and permute within quantile bins of this variable. [^11] This support-count measure captures a basic structural feature of the menu (how many distinct outcomes appear). It differs from the tradeoff complexity $\mathrm{TC}(A)$ used in Section 8, which isolates how much the two CDFs cross. We stratify on support count because it is the most direct determinant of rule activity patterns: rules that involve pairwise outcome comparisons (e.g., SAL, REG) mechanically depend on the number of support points. Concretely, for each rule $f$ independently and within each of $10$ quantile bins, we randomly reassign the $(L_{f}(A),A_{f}(A))$ pairs across menus within that bin. This preserves, for each rule and each complexity stratum, the marginal activity rate and recommendation rate, but destroys the mapping from a specific menu to its rule indicators.
The gate architecture, training protocol (50 folds, two-pass LR selection, same number of gradient steps and batch size as the main rule-gating model), and all other inputs-including the gate features $z(A)$-are identical. We repeat the permutation $50$ times and report the average test MSE across permutations.
@@ Results.
Table A.8 reports the comparison. The placebo library produces substantially higher MSE than the real library, confirming that the rule indicators carry genuine, menu-specific predictive content that cannot be replicated by indicators with matching marginal statistics. The gate’s predictive advantage is therefore attributable to the structured correspondence between menu features and rule recommendations, not merely to the availability of additional binary regressors.
Library Mean MSE SD $n$
Real (rule-gating) 0.01282 0.00059 50
Placebo (permuted) 0.02403 0.00032 50
Appendix M Feature-Based Baselines: Results
A natural question is whether competitive performance could be achieved by a standard off-the-shelf method applied to the same features, without the gate architecture or the behavioral structure of the RRM. To address this, we train two classical machine-learning models on the same data and evaluate them under the same 50-fold CV protocol as the main models.
Models.
The first is a fractional logit: a binomial generalized linear model (GLM) with a logit link, where the dependent variable is the continuous menu-level choice frequency $\hat{p}(A)\in[0,1]$ (following Papke and Wooldridge 1996). The $L_{2}$ regularization parameter adds a penalty proportional to $\lambda\|\beta\|_{2}^{2}$; $\lambda$ is selected by inner-fold validation from $\{0,0.01,0.1,1\}$. The second is a gradient-boosted tree ensemble (GBT, scikit-learn), a flexible nonparametric learner; the number of boosting rounds is selected from $\{100,200,500\}$ by inner-fold validation.
Feature sets.
Each model is trained on three alternative feature sets: (F1) summary statistics ($\approx$18 dimensions); (F2) rule indicators (24 dimensions); and (F3) the concatenation of F1 and F2 ($\approx$42 dimensions). Full details are in Appendix K.
Results.
Table A.9 reports results. On rule features alone (F2), rule-gating (MSE $0.01282$) substantially outperforms both the fractional logit ($0.02346$) and GBT ($0.02039$). The comparison does not isolate architecture from features (rule-gating also conditions on $z(A)$), but it shows that the activity-conditioned softmax gate extracts substantially more predictive content from the rule library than standard off-the-shelf methods.
GBT on summary statistics (F1, MSE $0.00997$) and on combined features (F3, MSE $0.00957$) outperform all structured models by an important margin. This shows that rich predictive information exists in the raw lottery features, and that a sufficiently flexible learner can exploit it. However, GBT produces an opaque mapping with no behavioral decomposition. Rule-gating therefore occupies a distinct niche: among all models that produce a transparent behavioral decomposition, it delivers the strongest prediction.
@@
As a robustness check, we estimate a variant in which the softmax gate operates on the raw menu encoding $\psi(A)\in\mathbb{R}^{4M}$ (Equation (21)) instead of the interpretable features $z(A)$. With $M=10$, this is a $40$-dimensional input, yielding $40\times|\mathcal{F}|$ gate parameters (plus biases) instead of $12\times|\mathcal{F}|$. Because the raw encoding is essentially injective on the menu support, the global identification conditions (G1)-(G2) of Theorem 1 are not applicable to this parameterization (see Section 6.6).
The raw-encoding gate achieves mean test MSE $0.01464$, compared to $0.01282$ for the baseline specification using $z(A)$ - an approximately $14\%$ increase (Figure A.5). The $12$ interpretable summary statistics therefore capture the decision-relevant variation at least as well as the raw encoding, while the additional dimensions of $\psi(A)$ appear to introduce noise. With more parameters to estimate, the raw-encoding gate is also more prone to overfitting.
@@
The global identification conditions (G1)–(G2) of Theorems 1 and 2 require that multiple menus map to the same feature value, so that within-feature variation in decisive-side patterns can identify the rule weights. When the gate operates on the raw menu encoding $\psi(A)\in\mathbb{R}^{4M}$, which is essentially injective on the menu support, this replication condition cannot hold: each menu maps to a distinct feature value, leaving (G1) vacuous.
For such high-dimensional encodings, identification can still be assessed through the local approach of Theorem 3. At a given parameter value $\vartheta^{\star}$, one computes the Jacobian $J(\vartheta^{\star})$ of the identification map in closed form (Section A.9) and checks whether it achieves full column rank. If so, part (a) of Theorem 3 guarantees that $\vartheta^{\star}$ is locally identified. The effective dimension of the identified parameter space can be read off from the rank: if $\mathrm{rank}(J)=r$, then $r$ independent parameter combinations are pinned down by the data, while the remaining KrK-r directions lie in the tangent null space.
This local approach complements the global conditions in the following sense. The coarsened features $z(A)$ used in the baseline specification are designed to satisfy (G1)–(G2), providing global identification guarantees that do not depend on a particular parameter value. When a researcher instead uses a richer feature map for which (G1) is not applicable, the Jacobian rank diagnostic provides a parameter-value-specific check that can be evaluated at the fitted parameters. In our empirical analysis (Section 9.1), both diagnostics support identification: the global conditions (G1)–(G2) hold for the $z(A)$ specification, and the Jacobian achieves full column rank at the fitted parameters for both the $z(A)$ and raw-$\psi$ specifications.

### 19 Mar 2026 v2

@@ Appendix C Proof of Corollary 2 (detailed version)
By (G1), $\mathrm{rank}(H^{(k)})=|\mathcal{F}|-1$, so the nullspace of $H^{(k)}$ is one-dimensional: there exists a nonzero vector $\bar{\omega}^{(k)}$ such that every solution to $H^{(k)}\omega=0$ is of the form $\omega=c\,\bar{\omega}^{(k)}$ for some scalar $c$. The true weight vector $\omega(x^{(k)})$ is strictly positive and lies in this nullspace, so $\bar{\omega}^{(k)}$ may be oriented to be strictly positive (set $\bar{\omega}^{(k)}:=\omega(x^{(k)})$). Every admissible solution—i.e., every $\omega\in\mathbb{R}_{++}^{|\mathcal{F}|}$ with $H^{(k)}\omega=0$—is then exactly a positive multiple $c\,\bar{\omega}^{(k)}$ with $c>0$. Thus, the menus with $\psi(A)=x^{(k)}$ identify $\omega(x^{(k)})$ up to a positive multiplicative scale factor.
@@
Appendix D Proof of Theorem 3
Proof.
Let $J:=|\mathcal{F}|$, and let the finite menu support be
$\mathcal{M}=\{A_{1},\ldots,A_{M}\}.$
Write
$p_{m}:=p(A_{m}),\qquad r_{m}:=\frac{p_{m}}{1-p_{m}},\qquad m=1,\ldots,M.$
For each menu $A_{m}$, define the sample share
$\hat{p}_{m}:=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}Y_{mi},$
and let $\hat{r}_{m}=\hat{p}_{m}/(1-\hat{p}_{m})$, with the understanding that any trimming to $\varepsilon_{n},1-\varepsilon_{n}]$ is asymptotically inactive under (E1) because $\varepsilon_{n}\downarrow 0$.
For each identifying feature value $x^{(k)}$, let
$\mathcal{M}_{k}=\{m_{k1},\ldots,m_{kM_{k}}\}$
denote the menus in the $k$th cell, i.e. the menus satisfying $\psi(A_{m})=x^{(k)}$. Define
$\omega^{(k)}:=\omega(x^{(k)}),\qquad\omega^{(k)}=(1,\eta^{(k)\top})^{\top},$
where the normalization $\omega_{f_{0}}(x^{(k)})=1$ is imposed for every $k$.
We prove the three claims in turn.
Steps 1–2: Joint asymptotic normality of odds ratios.
[…]
$\sqrt{N}(\hat{r}-r)\xrightarrow{d}\mathcal{N}(0,\Sigma_{r}),\qquad\Sigma_{r}=\operatorname{diag}\!\Big(\frac{p_{m}}{\pi_{m}(1-p_{m})^{3}}\Big).$ (A.16)
Trimming to $\varepsilon_{n},1-\varepsilon_{n}]$ is asymptotically inactive under (E1).
Step 3: From odds ratios to the restriction matrix $H^{(k)}$.
Fix a cell $k$. Recall that the $(m,f)$ entry of $H^{(k)}$ is $h_{f}(A_{m})=\kappa_{f}^{L}(A_{m})-r_{m}\,\kappa_{f}^{R}(A_{m})$. The decisive-side indicators $\kappa_{f}^{L},\kappa_{f}^{R}$ are deterministic functions of the menu, so $H^{(k)}$ depends on the data only through the odds ratios $r^{(k)}=(r_{k1},\ldots,r_{kM_{k}})^{\top}$, and does so affinely: there exists a deterministic matrix $B_{k}$ (constructed from the stacked indicators) such that $\mathrm{vec}(\hat{H}^{(k)}-H^{(k)})=B_{k}(\hat{r}^{(k)}-r^{(k)})$. Because this map is linear, the CLT for the odds ratios (A.16) carries over directly:
$\sqrt{N}\,\mathrm{vec}(\hat{H}^{(k)}-H^{(k)})\xrightarrow{d}\mathcal{N}(0,\Sigma_{H}^{(k)}),\qquad\Sigma_{H}^{(k)}:=B_{k}\,\Sigma_{r}^{(k)}\,B_{k}^{\top}.$
Step 4: Cellwise weight recovery.
Partition $H^{(k)}=[h_{0}^{(k)}\ \ H_{-}^{(k)}]$ where the first column corresponds to the baseline rule $f_{0}$. The normalization $\omega_{f_{0}}=1$ and the linear restriction $H^{(k)}\omega^{(k)}=0$ together imply $H_{-}^{(k)}\eta^{(k)}=-h_{0}^{(k)}$, where $\eta^{(k)}=(\omega_{f}(x^{(k)}))_{f\neq f_{0}}$ collects the non-baseline weights. Under (E2), $\operatorname{rank}H^{(k)}=J-1$, so the null space of $H^{(k)}$ is one-dimensional (by Theorem 1). This implies $H_{-}^{(k)}$ has full column rank J1J-1: if it did not, $\ker H^{(k)}$ would contain a vector with $\omega_{f_{0}}=0$ in addition to the true $\omega^{(k)}$, giving dimension $\geq 2$. The least-squares map
$\phi_{k}(H):=-(H_{-}^{\top}H_{-})^{-1}H_{-}^{\top}h_{0}$
is therefore well-defined and smooth in a neighborhood of $H^{(k)}$ (since the Gram matrix $H_{-}^{\top}H_{-}$ is nonsingular and matrix inversion is smooth). The delta method applied to $\hat{\eta}^{(k)}=\phi_{k}(\hat{H}^{(k)})$ yields
$\sqrt{N}\big(\hat{\eta}^{(k)}-\eta^{(k)}\big)\xrightarrow{d}\mathcal{N}\big(0,V_{\eta}^{(k)}\big),\qquad V_{\eta}^{(k)}=D\phi_{k}(H^{(k)})\,\Sigma_{H}^{(k)}\,D\phi_{k}(H^{(k)})^{\top},$
where $D\phi_{k}$ denotes the Jacobian of $\phi_{k}$ with respect to $\mathrm{vec}(H)$. Embedding via $\hat{\omega}^{(k)}=(1,\hat{\eta}^{(k)\top})^{\top}$ gives $\sqrt{N}(\hat{\omega}^{(k)}-\omega^{(k)})\xrightarrow{d}\mathcal{N}(0,V_{\omega}^{(k)})$, where $V_{\omega}^{(k)}$ is a $J\times J$ matrix obtained from $V_{\eta}^{(k)}$ by inserting a zero row and column in the $f_{0}$ position. This reflects the normalization $\omega_{f_{0}}\equiv 1$: the baseline weight is not estimated but fixed, so $\hat{\omega}_{f_{0}}^{(k)}-\omega_{f_{0}}^{(k)}=0$ with probability one and the $f_{0}$ coordinate contributes zero variance.
Step 5: Log-weight transformation and stacking across cells.
For each non-baseline rule $f$ and cell $k$, define $\hat{y}_{f}^{(k)}:=\log\hat{\omega}_{f}(x^{(k)})$ and $y_{f}^{(k)}:=\log\omega_{f}(x^{(k)})$. Assumption (E4) ensures $\omega_{f}(x^{(k)})\geq\underline{\omega}>0$, so the logarithm is smooth at the truth and the delta method gives
$\sqrt{N}\big(\hat{y}_{f}^{(k)}-y_{f}^{(k)}\big)\xrightarrow{d}\mathcal{N}\big(0,V_{y,f}^{(k)}\big),\qquad V_{y,f}^{(k)}=\omega_{f}(x^{(k)})^{-2}\,e_{f}^{\top}V_{\omega}^{(k)}e_{f}.$
To see this, note that $\hat{y}_{f}^{(k)}=g(\hat{\omega}^{(k)})$ where $g(\omega)=\log\omega_{f}$ depends only on the $f$-th coordinate. The gradient is $\nabla g(\omega)=e_{f}/\omega_{f}$, where $e_{f}$ is the $f$-th standard basis vector. The delta method gives asymptotic variance $\nabla g^{\top}V_{\omega}^{(k)}\nabla g=\omega_{f}(x^{(k)})^{-2}\,e_{f}^{\top}V_{\omega}^{(k)}e_{f}$. The quadratic form $e_{f}^{\top}V_{\omega}^{(k)}e_{f}$ is simply the $(f,f)$ diagonal entry of $V_{\omega}^{(k)}$, i.e. the asymptotic variance of the $f$-th weight estimator $\hat{\omega}_{f}^{(k)}$ in that cell.
Stacking across cells $k=1,\ldots,K$, the joint CLT for $\hat{y}_{f}:=(\hat{y}_{f}^{(1)},\ldots,\hat{y}_{f}^{(K)})^{\top}$ is
$\sqrt{N}\big(\hat{y}_{f}-y_{f}\big)\xrightarrow{d}\mathcal{N}(0,V_{y,f}),\qquad V_{y,f}=\operatorname{diag}(V_{y,f}^{(1)},\ldots,V_{y,f}^{(K)}).$
The diagonal structure holds because the cells partition the menus into disjoint groups with independent within-menu samples, so the first-stage estimators from distinct cells are asymptotically independent. This completes part (i).
Step 6: Second-stage asymptotics for the affine parameters.
Fix $f\neq f_{0}$. By construction $y_{f}=X_{K}\gamma_{f}$ where $\gamma_{f}=(\alpha_{f},\beta_{f}^{\top})^{\top}$ and $X_{K}$ stacks the augmented cell centroids $z_{k}=(1,x^{(k)\top})^{\top}$. Let $A_{f}:=(X_{K}^{\top}W_{f}X_{K})^{-1}X_{K}^{\top}W_{f}$, which is nonsingular by (E3). Then $\hat{\gamma}_{f}-\gamma_{f}=A_{f}(\hat{y}_{f}-y_{f})$, so
$\sqrt{N}(\hat{\gamma}_{f}-\gamma_{f})\xrightarrow{d}\mathcal{N}\big(0,V_{\gamma,f}\big),\qquad V_{\gamma,f}=A_{f}\,V_{y,f}\,A_{f}^{\top}.$
This proves part (ii).
Step 7: Functionals of the gate parameters.
Let $\theta=\{\gamma_{f}\}_{f\neq f_{0}}$ be the stacked parameter vector. By part (ii), $\sqrt{N}(\hat{\theta}-\theta_{0})\xrightarrow{d}\mathcal{N}(0,V_{\theta})$. For any continuously differentiable functional $\tau$, the delta method gives $\sqrt{N}(\tau(\hat{\theta})-\tau(\theta_{0}))\xrightarrow{d}\mathcal{N}(0,\nabla\tau(\theta_{0})^{\top}V_{\theta}\nabla\tau(\theta_{0}))$, proving part (iii). ∎
Appendix E Attention-Rule Removal
@@
Figures A.1 and A.2 replicate the complexity and risk mechanism plots under the 10-rule specification. The qualitative patterns persist: removing A1/A2 reallocates mass to the structured rules, but directional movements are unchanged.
@@
Appendix F Latent Gate vs. Effective Responsibility Weights
Figures A.3 and A.4 compare the effective (CA-normalized) responsibility weights $\tilde{q}_{f}(A;\theta)$ with the latent gate outputs $q_{f}(A;\theta)$ across deciles of tradeoff complexity and risk asymmetry. For each rule, the top panel shows the effective weight (as plotted in Figures 56) and the bottom panel shows the corresponding latent gate propensity before conditioning on activity. The qualitative patterns are preserved: for example, MMx rises with complexity and MAP declines under both measures. This confirms that the comparative-statics patterns in the main text reflect genuine variation in the learned gate, not mechanical artifacts of the activity normalization.
Figure A.3: Effective responsibility weight $\tilde{q}_{f}$ (top, blue) vs. latent gate output $q_{f}$ (bottom, orange) across deciles of tradeoff complexity, for the six highest-weight rules.
Figure A.4: Effective responsibility weight $\tilde{q}_{f}$ (top, blue) vs. latent gate output $q_{f}$ (bottom, orange) across deciles of risk asymmetry, for the six highest-weight rules.
Appendix G Paired Comparisons
@@
Context-dependent Rule-gating test_mse -0.00045 0.00010 50
Neural CPT Rule-gating test_mse 0.00756 0.00014 50
Neural EU Rule-gating test_mse 0.00826 0.00012 50
Neural PT Rule-gating test_mse 0.00675 0.00013 50
Rule-gating Rule-gating test_mse -0.00221 0.00006 50
Value-based Rule-gating test_mse 0.00532 0.00012 50
MOT Rule-gating test_mse -0.00250 0.00012 50
Appendix H Model-Level Uncertainty Intervals
@@
Appendix I Ablation Effect Split-Variability Intervals
@@
The leading ablation effects are SAL2 and SAL, followed by REGmed and REG. At the other extreme, MMa has a small negative ablation effect (removing it slightly improves MSE), indicating that its contribution is fully substitutable by the remaining rules.
@@
Appendix J Decomposition Stability
@@
A1 0.2882 0.0098 0.03
A2 0.2631 0.0092 0.03
SAL 0.0927 0.0060 0.06 0.0471 0.0271 0.57
REG 0.0717 0.0056 0.08 0.0123 0.0270 2.19
MAP 0.0595 0.0047 0.08 0.0119 0.0193 1.62
MMx 0.0500 0.0058 0.12 0.0118 0.0263 2.23
MMn 0.0432 0.0143 0.33 0.0095 0.0177 1.86
SAL2 0.0385 0.0022 0.06 0.0529 0.0300 0.57
REGmed 0.0373 0.0085 0.23 0.0178 0.0210 1.18
DISmed 0.0268 0.0076 0.28 0.0005 0.0179 39.76
DIS 0.0218 0.0104 0.48 0.0041 0.0174 4.28
MMa 0.0073 0.0048 0.65 -0.0032 0.0157 4.98
Note: Spearman rank correlation of $w_{f}$ orderings across all fold pairs: mean $\rho=0.930$, range $[0.762,1.000]$. Spearman rank correlation of $\phi(f)$ orderings: mean $\rho=0.442$, range [0.636,0.952][-0.636,0.952].
Appendix K Selection Stability
@@
Appendix L Feature-Based Baselines
We train two baseline models—a fractional logit (Papke and Wooldridge, 1996) and a gradient-boosted tree ensemble (GBT, scikit-learn)—on three feature sets: (F1) summary statistics ($\approx$18 dim), (F2) rule indicators (24 dim), and (F3) concatenation ($\approx$42 dim). All use the same 50-fold CV protocol. The fractional logit is a binomial GLM with logit link on the continuous choice frequency $\hat{p}(A)\in[0,1]$; $L_{2}$ regularization is selected from $\{0,0.01,0.1,1\}$ by inner-fold validation. GBT uses max_depth=4 with the number of estimators selected from $\{100,200,500\}$.
@@ Results.
Table A.7 reports results. On rule features alone (F2), rule-gating substantially outperforms both baselines. This comparison isolates the gating mechanism’s contribution (not a pure feature comparison, since rule-gating also uses $z(A)$). GBT on combined features (F3) outperforms all structured models but provides no behavioral decomposition.
@@
Appendix M Placebo Rule Library
We construct a placebo test by randomly reassigning each rule’s $(L_{f}(A),A_{f}(A))$ pairs across menus within $10$ complexity strata, [^14] Stratified by support-count quantiles to preserve marginal activity rates per stratum. preserving marginal statistics but destroying menu-specific content ($50$ repetitions). Table A.8 shows that the placebo library produces substantially higher MSE, confirming that the rule indicators carry genuine, menu-specific predictive content.
Library Mean MSE SD $n$
Real (rule-gating) 0.01168 0.00049 50
Placebo (permuted) 0.02197 0.00032 50
@@
We estimate a variant using the raw encoding $\psi(A)\in\mathbb{R}^{40}$ instead of $z(A)$, yielding $40\times|\mathcal{F}|$ gate parameters. The raw-encoding gate achieves MSE $0.0139$ vs. $0.0117$ for the baseline—about $19\%$ worse (Figure A.5)—indicating that the interpretable summary statistics capture decision-relevant variation without the overfitting induced by higher-dimensional inputs.
@@
The global conditions (G1)–(G2) require within-feature replication, which fails when the gate operates on the essentially injective raw encoding $\psi(A)\in\mathbb{R}^{4M}$. For such specifications, identification is assessed locally via Theorem A.2: if the Jacobian $J(\vartheta^{\star})$ of the identification map (Section A.9) has full column rank at the fitted parameters, part (a) guarantees local identification.
In Section 7.1, both diagnostics support identification: the global conditions hold for the $z(A)$ specification, and the Jacobian achieves full column rank at the fitted parameters for both the $z(A)$ and raw-$\psi$ specifications, where “full column rank” is evaluated relative to the effective parameter dimension (accounting for the redundancy in $z(A)$ that reduces its dimension from $12$ to $d_{\mathrm{eff}}=11$).